OpenStudy (anonymous):

If f(x) = 2x − 4 and g(x) = x2 − 4x, find f(3)-1.

OpenStudy (shamim):

\[f(x)=2x-4\]

OpenStudy (shamim):

\[f(3)=2 \times 3-4=2\]

OpenStudy (shamim):

\[f(3)-1=2-1=1\]